### v2

# Estimators for multivariate allometric regression model

Abstract
In a regression model with multiple response variables and multiple explanatory variables, if the difference of the mean vectors of the response variables for different values of explanatory variables is always in the direction of the first principal eigenvector of the covariance matrix of the response variables, then it is called a multivariate allometric regression model. This paper studies the estimation of the first principal eigenvector in the multivariate allometric regression model. A class of estimators that includes conventional estimators is proposed based on weighted sum-of-squares matrices of regression sum-of-squares matrix and residual sum-of-squares matrix. We establish an upper bound of the mean squared error of the estimators contained in this class, and the weight value minimizing the upper bound is derived. Sufficient conditions for the consistency of the estimators are discussed in weak identifiability regimes under which the difference of the largest and second largest eigenvalues of the covariance matrix decays asymptotically and in “large $p$, large $n$” regimes, where $p$ is the number of response variables and $n$ is the sample size. Several numerical results are also presented.
Keywords: covariance matrix, multivariate regression model, principal component analysis, sum-of-squares matrix, weak identifiability.
Mathematics Subject Classification: 62J05, 62H12, 62H25
1 Introduction
The multivariate regression model, a regression model with multiple response variables and multiple explanatory variables, is one of the most popular models in multivariate analysis; see Anderson (2003), Johnson and Wichern (2007), Rencher and Christensen (2012), and so on. For the model, the method of least squares works well. When additional assumptions such as low rankness of the coefficient matrix exist, there is another estimator that is generally better than the ordinary least squares estimator. In particular, we consider the multivariate allometric regression model, proposed in Tarpey and Ivey (2006), which is a reduced rank regression model of rank one with the constraint that the differences of the mean vectors of response variables for different values of explanatory variables are in the direction of the first principal eigenvector of the covariance matrix. The first principal eigenvector is a key quantity to infer the multivariate allometric regression model, so we consider the estimation of this quantity. Specifically, we define a class of estimators that includes natural conventional estimators, evaluate the mean squared errors of the estimators contained in this class, compare them from the viewpoint of the consistency, propose a new estimator, and examine the proposed estimator through numerical simulations.
Let
$\bm{Y}=\bm{1}_{n}\bm{\mu}^{\prime}+\bm{X}\bm{B}+\bm{E}\ \ \ {\rm with}\ \mathsf{E}[{\rm vec}(\bm{E})]=\bm{0}_{np},\ \mathsf{V}[{\rm vec}(\bm{E})]=\bm{\Sigma}\otimes\bm{I}_{n}$ ( 1.1)
be a multivariate regression model, where $\bm{Y}$ is an $n\times p$ matrix of response variables ($n$ is the sample size and $p$ is the number of response variables), $\bm{1}_{n}$ is the $n\times 1$ vector whose all elements are $1$, $\bm{\mu}$ is a $p\times 1$ vector of intercepts, $\bm{X}$ is an $n\times q$ matrix of explanatory variables ($q$ is the number of explanatory variables) such that each explanatory variable is standardized to have mean 00 (i.e., $\bm{X}^{\prime}\bm{1}_{n}=\bm{0}_{q}$), $\bm{B}$ is a $q\times p$ matrix of regression coefficients, $\bm{E}$ is an $n\times p$ matrix of errors with mean 00, $\bm{\Sigma}$ is a $p\times p$ positive-definite matrix denoting the common covariance matrix of errors of the $p$ response variables, $\bm{0}_{np}$ is the $np\times 1$ zero vector, $\bm{I}_{n}$ is the $n\times n$ identity matrix, and $\otimes$ denotes the Kronecker product. Putting $\bm{Y}=\left(\bm{y}_{1},\bm{y}_{2},\dots,\bm{y}_{n}\right)^{\prime}$, $\bm{X}=\left(\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{n}\right)^{\prime}$, and $\bm{E}=\left(\bm{e}_{1},\bm{e}_{2},\dots,\bm{e}_{n}\right)^{\prime}$, we can rewrite (1.1) as
$\bm{y}_{i}=\bm{\mu}+\bm{B}^{\prime}\bm{x}_{i}+\bm{e}_{i}\ \ \ {\rm with}\ \mathsf{E}[\bm{e}_{i}]=\bm{0}_{p},\ \mathsf{V}[\bm{e}_{i}]=\bm{\Sigma},\ \ \ i=1,2,\dots,n.$
Each response vector $\bm{y}_{i}$ has the mean vector $\bm{\mu}+\bm{B}^{\prime}\bm{x}_{i}$ and the covariance matrix $\bm{\Sigma}$. The multivariate allometric regression model is a special case of the multivariate regression model. Let $\lambda_{1},\lambda_{2},\dots,\lambda_{p}$ denote the ordered eigenvalues of $\bm{\Sigma}$ ($\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{p}>0$), and let $\bm{\gamma}_{i}$ be a normalized principal eigenvector of $\bm{\Sigma}$ corresponding to its eigenvalue $\lambda_{i}$ for $i=1,\ldots,p$. Suppose that $\lambda_{1}>\lambda_{2}$ holds, which implies that $\bm{\gamma}_{1}$ is unique up to a sign change. Then, the multivariate allometric regression model is expressed as
$\bm{Y}=\bm{1}_{n}\bm{\mu}^{\prime}+\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{E}\ \ \ {\rm with}\ \mathsf{E}[{\rm vec}(\bm{E})]=\bm{0}_{np},\ \mathsf{V}[{\rm vec}(\bm{E})]=\bm{\Sigma}\otimes\bm{I}_{n}=\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}\otimes\bm{I}_{n},$ ( 1.2)
where $\bm{\alpha}$ is a $q\times 1$ vector, $\bm{\Gamma}=\left(\bm{\gamma}_{1},\bm{\gamma}_{2},\dots,\bm{\gamma}_{p}\right)$ is a $p\times p$ orthogonal matrix, and $\bm{\Lambda}={\rm diag}\left(\lambda_{1},\lambda_{2},\dots,\lambda_{p}\right)$ is a $p\times p$ diagonal matrix. We assume the multivariate normality of the distribution of the error matrix $\bm{E}$: ${\rm vec}(\bm{E})\sim\mathcal{N}_{np}\left(\bm{0}_{np},\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}\otimes\bm{I}_{n}\right)$. In this setup, we consider the estimation of $\bm{\gamma}_{1}$ in (1.2).
The multivariate allometric regression model was proposed in Tarpey and Ivey (2006) as an extension of the allometric extension model (Flury, 1997; Hills, 2006) to the context of multivariate regression. The allometric extension model means that several population distributions share a first principal eigenvector of their covariance matrices and the differences of their mean vectors are in the direction of the first principal eigenvector, which has been applied to various areas including biometrics (Klingenberg and Froese, 1991; Flury, 1997). Bartoletti et al. (1999) proposed a hypothesis testing procedure for the allometric extension model based on a $\chi^{2}$ statistic in large sample situations. Kurata et al. (2008) derived conditions for multivariate conditional distributions to form the allometric extension model. Matsuura and Kurata (2014) discussed principal points of multivariate mixture distributions of the allometric extension model. Tsukuda and Matsuura (2023) proposed hypothesis testing procedures for the allometric extension model in high-dimensional and strongly spiked eigenvalue situations. Kawamoto et al. (2023) evaluated the performance of spectral clustering algorithm when it is used for a random sample from the mixture of two subpopulations that are allometric extensions of each other.
The multivariate allometric regression model can be viewed as a partial generalization of the allometric extension model, because the multivariate allometric regression model (1.2) can deal with infinite mixture of population distributions incorporating multivariate regression models with explanatory variables, while there is an restriction that each population distribution has a common covariance matrix of errors. Hypothesis testing for the multivariate allometric regression model has been discussed in Schott (2017). The multivariate allometric regression model is also closely related to multivariate reduced rank regression (Anderson, 1951, 2003; Izenman, 1975; Reinsel et al., 2022). If the differences of the mean vectors of the response variables for different explanatory variables are in a common direction but do not necessarily coincide with the first principal eigenvector $\bm{\gamma}_{1}$ of $\bm{\Sigma}$, then (1.2) reduces to the multivariate reduced rank regression of rank one. However, in (1.2), using the information that the common direction of the differences of the mean vectors coincides with the first principal eigenvector may enable more accurate estimation compared with the multivariate reduced rank regression of rank one.
This paper studies the estimation of $\bm{\gamma}_{1}$ in (1.2). From the viewpoint of principal component analysis, inferring the first principal eigenvector $\bm{\gamma}_{1}$ is of special interest. Moreover, although we focus on the estimation of $\bm{\gamma}_{1}$, precise estimation of $\bm{\gamma}_{1}$ may lead to precise estimation of $\bm{B}=\bm{\alpha}\bm{\gamma}_{1}^{\prime}$. In fact, once an estimator $\hat{\bm{\gamma}}_{1}$ of $\bm{\gamma}_{1}$ is obtained, an estimator of $\bm{B}$ is obtained as $\hat{\bm{B}}(\hat{\bm{\gamma}}_{1}\hat{\bm{\gamma}}_{1}^{\prime})=\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{Y}\hat{\bm{\gamma}}_{1}\hat{\bm{\gamma}}_{1}^{\prime}$ if $\bm{X}^{\prime}\bm{X}$ is invertible. In the reduced rank regression, such estimators are generally better than the ordinary least squares estimator $\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{Y}$; see Anderson (1999). Also when (1.2) is supposed, $\hat{\bm{B}}(\hat{\bm{\gamma}}_{1}\hat{\bm{\gamma}}_{1}^{\prime})$ is considered to be better than the ordinary least squares estimator in general. This is another motivation to focus on estimating $\bm{\gamma}_{1}$. In particular, the consistency up to sign of $\hat{\bm{\gamma}}_{1}$ is important. That is because, e.g., under some adequate asymptotic regimes, if $\hat{\bm{\gamma}}_{1}$ is consistent, then $\hat{\bm{B}}(\hat{\bm{\gamma}}_{1}\hat{\bm{\gamma}}_{1}^{\prime})$ is also consistent.
In this paper, we deal with a class of estimators of $\bm{\gamma}_{1}$ based on weighted matrices of two independent sum-of-squares matrices: regression sum-of-squares matrix and residual sum-of-squares matrix. Non-asymptotic upper bounds are established for the mean squared errors of the estimators, and the weight value minimizing the upper bound for the mean squared error is derived. Subsequently, we discuss the consistency of the estimators under weak identifiability regimes under which $\lambda_{1}-\lambda_{2}\to 0$ as $n\to\infty$, and “large $p$, large $n$” regimes under which $p$ and $n$ tend to $\infty$ simultaneously. As remarked in Paindaveine et al. (2020), properties of principal component analysis under weak identifiability regimes have not been sufficiently investigated. Under large $p$, large $n$ regimes that include so-called high-dimensional asymptotic regimes $n,p\to\infty$ with $p/n\to\infty$, we consider weak identifiability regimes and a sort of spiked eigenvalue regimes simultaneously. In the context of multivariate regression, large $p$, large $n$ regimes have been discussed in a lot of works such as Bing and Wegkamp (2019), Bunea et al. (2011), Chen et al. (2013), Giraud (2011), She and Chen (2017), Yanagihara et al. (2015), Yuan et al. (2007), and so on. Furthermore, numerical results are presented to compare the proposed estimator with some conventional estimators.
2 Conventional estimators
This section presents some conventional estimators of $\bm{\gamma}_{1}$ in the multivariate allometric regression model (1.2). Here and subsequently, assume that $n>1+q$ and $\mathrm{rank}(\bm{X})=q$. Note that $n>2+q$ will be assumed in Section 6 in order to avoid zero in a denominator. Let
$\bm{S}_{R}=\bm{Y}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{Y},$
$\bm{S}_{E}=\bm{Y}^{\prime}\left\{\bm{I}_{n}-\frac{1}{n}\bm{1}_{n}\bm{1}_{n}^{\prime}-\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\right\}\bm{Y},$
$\bm{S}_{T}=\bm{Y}^{\prime}\left(\bm{I}_{n}-\frac{1}{n}\bm{1}_{n}\bm{1}_{n}^{\prime}\right)\bm{Y}.$
The matrices $\bm{S}_{R}$, $\bm{S}_{E}$, and $\bm{S}_{T}$ are so-called regression sum-of-squares, residual sum-of-squares, and total sum-of-squares matrices, respectively. It is obvious that $\bm{S}_{T}=\bm{S}_{R}+\bm{S}_{E}$. Straightforward calculations give
$\mathsf{E}\left[\bm{S}_{R}\right]=q\bm{\Sigma}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}=q\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}+\|\bm{X}\bm{\alpha}\|^{2}\bm{\gamma}_{1}\bm{\gamma}_{1}^{\prime},$
$\mathsf{E}\left[\bm{S}_{E}\right]=(n-1-q)\bm{\Sigma}=(n-1-q)\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime},$
$\mathsf{E}\left[\bm{S}_{T}\right]=(n-1)\bm{\Sigma}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}=(n-1)\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}+\|\bm{X}\bm{\alpha}\|^{2}\bm{\gamma}_{1}\bm{\gamma}_{1}^{\prime}.$
Let $\hat{\bm{\gamma}}_{1}^{(R)}$, $\hat{\bm{\gamma}}_{1}^{(E)}$, and $\hat{\bm{\gamma}}_{1}^{(T)}$ be the first principal eigenvectors of $\bm{S}_{R}$, $\bm{S}_{E}$, and $\bm{S}_{T}$, respectively, which estimate $\bm{\gamma}_{1}$. The estimator $\hat{\bm{\gamma}}_{1}^{(R)}$ is a typical estimator to estimate the regression coefficient matrix under reduced rank regression of rank one; see Reinsel et al. (2022). However, in the multivariate allometric regression model (1.2), $\hat{\bm{\gamma}}_{1}^{(R)}$ does not sufficiently use the information that the first principal eigenvector of $\bm{\Sigma}$ is $\bm{\gamma}_{1}$. The estimator $\hat{\bm{\gamma}}_{1}^{(E)}$ only uses this information, not using the information that $\bm{B}$ is decomposed as $\bm{B}=\bm{\alpha}\bm{\gamma}_{1}^{\prime}$. The estimator $\hat{\bm{\gamma}}_{1}^{(T)}$ has been presented in Tarpey and Ivey (2006), which uses both the informations.
However, it is natural to consider that $\hat{\bm{\gamma}}^{(T)}$ is not always the best estimator. To be more specific, for example, we consider the accuracy of the first principal eigenvector $\hat{\bm{\gamma}}^{(E)}_{1}$ of $\bm{S}_{E}$. As $n\to\infty$ with $q$ held fixed, under the assumption that $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a positive-definite matrix, $\sqrt{n}(\hat{\bm{\gamma}}^{(E)}_{1}-\bm{\gamma}_{1})$ converges in distribution to
$\mathcal{N}_{p}\left(\bm{0}_{p},\lambda_{1}\sum_{i=2}^{p}\frac{\lambda_{i}}{(\lambda_{1}-\lambda_{i})^{2}}\bm{\gamma}_{i}\bm{\gamma}_{i}^{\prime}\right);$
see, e.g., Kollo and Neudecker (1993). This suggests that if $\lambda_{1}-\lambda_{2}$ is small, then $\hat{\bm{\gamma}}^{(E)}_{1}$ is not a good estimator. Hence, we consider that this property may be inherited by $\hat{\bm{\gamma}}^{(T)}_{1}$, since $\bm{S}_{T}=\bm{S}_{R}+\bm{S}_{E}$ includes $\bm{S}_{E}$. This estimate will be numerically examined in Section 7.
Remark 1.
Note that $\hat{\bm{\gamma}}_{1}^{(T)}$ is the first principal eigenvector in the usual principal component analysis for the multivariate location-scale model that is given by $\bm{\alpha}=\bm{0}_{q}$ in (1.2).
Remark 2.
Paindaveine et al. (2020) studied consistency and asymptotic normality of sample first principal eigenvector under weak identifiability regimes with a specific form of the covariance matrix. Moreover, considering elliptical distributions, Paindaveine and Verdebout (2023) studied asymptotic properties of the leading eigenvector of the Tyler shape estimator.
Remark 3.
In our setup, $\hat{\bm{\gamma}}_{1}^{(E)}$ is the first principal eigenvector of a Wishart matrix with $n-q-1$ degrees of freedom and scale matrix parameter $\bm{\Sigma}$. Thus, if $q$ is fixed, properties of $\hat{\bm{\gamma}}_{1}^{(E)}$ can be interpreted as general properties for the Wishart distribution.
3 Estimators based on weighted matrices of $\bm{S}_{R}$ and $\bm{S}_{E}$
Let us introduce a class of estimators of $\bm{\gamma}_{1}$ based on weighted matrices of $\bm{S}_{R}$ and $\bm{S}_{E}$, and discuss some properties of the estimators such as their mean squared errors and consistency. Let $\hat{\bm{\gamma}}_{1}(w)$ be the first principal eigenvector of a weighted matrix of $\bm{S}_{R}$ and $\bm{S}_{E}$:
$\bm{S}(w)=(1-w)\bm{S}_{R}+w\bm{S}_{E}$
with $0\leq w\leq 1$. Obviously, $\hat{\bm{\gamma}}_{1}(0)$, $\hat{\bm{\gamma}}_{1}(0.5)$, and $\hat{\bm{\gamma}}_{1}(1)$ correspond to $\hat{\bm{\gamma}}_{1}^{(R)}$, $\hat{\bm{\gamma}}_{1}^{(T)}$, and $\hat{\bm{\gamma}}_{1}^{(E)}$, respectively. In this section, we establish an upper bound for the mean squared error of the estimator $\hat{\bm{\gamma}}_{1}(w)$.
As a preparation, we derive a simple form of the expectation of the squared Frobenius norm of $\bm{S}(w)-\mathsf{E}[\bm{S}(w)]$.
Lemma 3.1.
It holds that
$\mathsf{E}\left[\left\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right\|_{\rm F}^{2}\right]=\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+2(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}.$
Proof.
From the assumption of ${\rm vec}(\bm{E})\sim\mathcal{N}_{np}\left(\bm{0}_{np},\bm{\Sigma}\otimes\bm{I}_{n}\right)$, it follows that $\bm{S}_{R}$ and $\bm{S}_{E}$ are independent each other; see, e.g., Anderson(2003). On account of this independence, we have
$\mathsf{E}\left[\left\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right\|_{\rm F}^{2}\right]$
$=\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right)^{\prime}\left(\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right)\right)\right]=\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right)^{2}\right)\right]$
$=(1-w)^{2}\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}]\right)^{2}\right)\right]+w^{2}\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{E}-\mathsf{E}[\bm{S}_{E}]\right)^{2}\right)\right].$
First, we evaluate $\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}]\right)^{2}\right)\right]$ as
$\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}]\right)^{2}\right)\right]$
$=\mathsf{E}\left[{\rm tr}\left(\left\{\bm{Y}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{Y}-\left(q\bm{\Sigma}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\right)\right\}^{2}\right)\right]$
$=\mathsf{E}\left[{\rm tr}\left(\left\{\left(\bm{1}_{n}\bm{\mu}^{\prime}+\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{E}\right)^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\left(\bm{1}_{n}\bm{\mu}^{\prime}+\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{E}\right)-\left(q\bm{\Sigma}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\right)\right\}^{2}\right)\right]$
$=\mathsf{E}\left[{\rm tr}\left(\left\{\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}+\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}-q\bm{\Sigma}\right\}^{2}\right)\right].$
We note that $\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}$ is distributed as the $p$-dimensional Wishart distribution $\mathcal{W}_{p}\left(q,\bm{\Sigma}\right)$ with $q$ degrees of freedom and scale matrix parameter $\bm{\Sigma}$. It is well known that if $\bm{W}\sim\mathcal{W}_{p}\left(q,\bm{\Sigma}\right)$, then $\mathsf{E}\left[\bm{W}\right]=q\bm{\Sigma}$ and $\mathsf{E}\left[{\rm tr}\left(\bm{W}^{2}\right)\right]=(q+q^{2}){\rm tr}\left(\bm{\Sigma}^{2}\right)+q\left({\rm tr}\left(\bm{\Sigma}\right)\right)^{2}$. Moreover, as it will be seen in Lemma 3.2 below, it also holds that $\mathsf{E}\left[\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}\bm{E}^{\prime}\bm{X}\right]=\bm{0}_{p\times q}$. Hence, letting $\bm{W}=\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}$, we have
$\mathsf{E}\left[{\rm tr}\left(\left\{\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}+\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}-q\bm{\Sigma}\right\}^{2}\right)\right]$
$=\mathsf{E}\left[{\rm tr}\left(\left(\bm{W}+\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}-q\bm{\Sigma}\right)^{2}\right)\right]$
$=\mathsf{E}\Bigl{[}{\rm tr}\Bigl{(}\Bigl{(}\bm{W}^{2}-q\bm{W}\bm{\Sigma}-q\bm{\Sigma}\bm{W}+q^{2}\bm{\Sigma}^{2}$
$\quad+\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}+\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}+\bm{\gamma}_{1}\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}\bm{E}^{\prime}\bm{X}\bm{\alpha}\bm{\gamma}_{1}^{\prime}\Bigr{)}\Bigr{)}\Bigr{]}$
$=\mathsf{E}\Bigl{[}{\rm tr}\Bigl{(}\bm{W}^{2}-q^{2}\bm{\Sigma}^{2}\Bigr{)}+2\left(\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}\bm{\gamma}_{1}\right)^{2}+2\bm{\alpha}^{\prime}\bm{X}^{\prime}\bm{E}\bm{E}^{\prime}\bm{X}\bm{\alpha}\Bigr{]}$
$=q{\rm tr}(\bm{\Sigma}^{2})+q\left({\rm tr}(\bm{\Sigma})\right)^{2}+2\lambda_{1}\|\bm{X}\bm{\alpha}\|^{2}+2{\rm tr}(\bm{\Sigma})\|\bm{X}\bm{\alpha}\|^{2}.$
Next, we evaluate $\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{E}-\mathsf{E}[\bm{S}_{E}]\right)^{2}\right)\right]$. Since $\bm{S}_{E}\sim\mathcal{W}_{p}\left(n-1-q,\bm{\Sigma}\right)$, it holds that
$\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{E}-\mathsf{E}[\bm{S}_{E}]\right)^{2}\right)\right]$
$=\mathsf{E}\left[{\rm tr}\left(\bm{S}_{E}^{2}\right)\right]-{\rm tr}\left(\left(\mathsf{E}[\bm{S}_{E}]\right)^{2}\right)$
=$=\left\{(n-1-q)+(n-1-q)^{2}\right\}{\rm tr}\left(\bm{\Sigma}^{2}\right)+(n-1-q)\left({\rm tr}\left(\bm{\Sigma}\right)\right)^{2}-(n-1-q)^{2}{\rm tr}\left(\bm{\Sigma}^{2}\right)$
$=(n-1-q){\rm tr}(\bm{\Sigma}^{2})+(n-1-q)\left({\rm tr}(\bm{\Sigma})\right)^{2}.$
We see from the above calculations that
$\mathsf{E}\left[\left\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\right\|_{\rm F}^{2}\right]$
$=(1-w)^{2}\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}]\right)^{2}\right)\right]+w^{2}\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{E}-\mathsf{E}[\bm{S}_{E}]\right)^{2}\right)\right]$
$=(1-w)^{2}\left\{q{\rm tr}(\bm{\Sigma}^{2})+q\left({\rm tr}(\bm{\Sigma})\right)^{2}+2\left(\lambda_{1}+{\rm tr}(\bm{\Sigma})\right)\|\bm{X}\bm{\alpha}\|^{2}\right\}$
$\quad+w^{2}\left\{(n-1-q){\rm tr}(\bm{\Sigma}^{2})+(n-1-q)\left({\rm tr}(\bm{\Sigma})\right)^{2}\right\}$
=$=\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+2(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}.$
This completes the proof. ∎
The following lemma is already used in the proof above.
Lemma 3.2.
It holds that
$\mathsf{E}\left[\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}\bm{E}^{\prime}\bm{X}\right]=\bm{0}_{p\times q}.$
Proof.
Let $\bm{g}_{1}^{\prime},\bm{g}_{2}^{\prime},\dots,\bm{g}_{q}^{\prime}$ be the row vectors of $\bm{G}=\left(\bm{X}^{\prime}\bm{X}\right)^{-1/2}\bm{X}^{\prime}\bm{E}$, that is, $\bm{G}=\left(\bm{g}_{1},\bm{g}_{2},\dots,\bm{g}_{q}\right)^{\prime}$, where $(\bm{X}^{\prime}\bm{X})^{-1/2}$ is the $q\times q$ symmetric matrix satisfying that $\{(\bm{X}^{\prime}\bm{X})^{-1/2}\}^{2}=\left(\bm{X}^{\prime}\bm{X}\right)^{-1}$. Then, from $\mathsf{V}[{\rm vec}(\bm{G})]=\bm{\Sigma}\otimes\bm{I}_{q}$, it follows that $\bm{g}_{1},\bm{g}_{2},\dots,\bm{g}_{q}$ are independent and identically distributed as $\mathcal{N}_{p}(\bm{0}_{p},\bm{\Sigma})$, and thus we see that
$\mathsf{E}\left[\bm{E}^{\prime}\bm{X}\left(\bm{X}^{\prime}\bm{X}\right)^{-1}\bm{X}^{\prime}\bm{E}\bm{E}^{\prime}\bm{X}\right]=\mathsf{E}\left[\bm{G}^{\prime}\bm{G}\bm{G}^{\prime}\left(\bm{X}^{\prime}\bm{X}\right)^{1/2}\right]=\mathsf{E}\left[\sum_{i=1}^{q}\bm{g}_{i}\bm{g}_{i}^{\prime}\left(\bm{g}_{1},\bm{g}_{2},\dots,\bm{g}_{q}\right)\right]\left(\bm{X}^{\prime}\bm{X}\right)^{1/2}=\bm{0}_{p\times q}.$
This completes the proof. ∎
Now, we establish an upper bound, without any unspecified constant, for the mean squared error (up to sign) of the estimator $\hat{\bm{\gamma}}_{1}(w)$:
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|^{2}\right]$
in the following theorem.
Theorem 3.3.
It holds that
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|^{2}\right]$
$\leq\frac{8\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}.$ ( 3.1)
Proof.
First, let $\lambda_{1}(w)$ and $\lambda_{2}(w)$ be the largest and the second largest eigenvalues of $\mathsf{E}[\bm{S}(w)]$, respectively. It follows from
$\mathsf{E}[\bm{S}(w)]$ $=\mathsf{E}\left[(1-w)\bm{S}_{R}+w\bm{S}_{E}\right]$
$=(1-w)\left(q\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}+\|\bm{X}\bm{\alpha}\|^{2}\bm{\gamma}_{1}\bm{\gamma}_{1}^{\prime}\right)+w(n-1-q)\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}$
=$=\left\{q(1-w)+(n-1-q)w\right\}\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\bm{\gamma}_{1}\bm{\gamma}_{1}^{\prime}$
that
$\lambda_{1}(w)=\left\{q(1-w)+(n-1-q)w\right\}\lambda_{1}+(1-w)\|\bm{X}\bm{\alpha}\|^{2},$
$\lambda_{2}(w)=\left\{q(1-w)+(n-1-q)w\right\}\lambda_{2},$
and that $\bm{\gamma}_{1}$ is the principal eigenvector of $\mathsf{E}[\bm{S}(w)]$ corresponding to $\lambda_{1}(w)$. The Davis–Kahan theorem implies that
$\sin\angle(\bm{\gamma}_{1},\hat{\bm{\gamma}}_{1}(w))\leq\frac{2\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\|_{{\rm op}}}{\lambda_{1}(w)-\lambda_{2}(w)},$
where the angle $\angle$ is taken in $[0,\pi/2]$, $\|\cdot\|_{{\rm op}}$ denotes the operator norm; see, for example, Nadler (2008) and Vershynin (2018) for the Davis–Kahan theorem. Noting that $\left|\bm{\gamma}_{1}^{\prime}\hat{\bm{\gamma}}_{1}(w)\right|\leq 1$ and $\|\cdot\|_{{\rm op}}\leq\|\cdot\|_{\rm F}$, we have
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|^{2}\right]=\mathsf{E}\left[2-2\max_{\theta=-1,1}\left(\theta\bm{\gamma}_{1}^{\prime}\hat{\bm{\gamma}}_{1}(w)\right)\right]$
$\leq\mathsf{E}\left[2-2\left(\bm{\gamma}_{1}^{\prime}\hat{\bm{\gamma}}_{1}(w)\right)^{2}\right]=\mathsf{E}\left[2\left(\sin\angle(\bm{\gamma}_{1},\hat{\bm{\gamma}}_{1}(w))\right)^{2}\right]$
$\leq\mathsf{E}\left[2\frac{4\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\|_{{\rm op}}^{2}}{\left(\lambda_{1}(w)-\lambda_{2}(w)\right)^{2}}\right]$
$\leq\mathsf{E}\left[\frac{8\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\|_{\rm F}^{2}}{\left[\left\{q(1-w)+(n-1-q)w\right\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}\right]$
$=\frac{8\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}},$
which completes the proof. ∎
Remark 4.
Although the estimation of $\bm{\gamma}_{1}$ is focused on, corresponding results for other eigenvectors of $\bm{S}(w)$ follows by the almost same method under adequate assumptions on $\bm{\Sigma}$.
Remark 5.
When $\bm{\alpha}=\bm{0}_{q}$, the eigenvector $\hat{\bm{\gamma}}_{1}^{(T)}=\hat{\bm{\gamma}}(1/2)$ of the sample covariance matrix $\bm{S}_{T}$ corresponding to the largest eigenvalue satisfies that
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}^{(T)}-\bm{\gamma}_{1}\|^{2}\right]\leq\frac{8\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}}{(n-1)(\lambda_{1}-\lambda_{2})^{2}}.$
Using Theorem 3.3, we will discuss sufficient conditions for the consistency up to sign of $\hat{\bm{\gamma}}_{1}(w)$ under some asymptotic regimes, as shown in Sections 4 and 5. As mentioned in Remark 7, when the dimension $p$ is held fixed, the bound of Theorem 3.3 seems sharp from the viewpoint of consistency. Here the meaning of the consistency up to sign of $\hat{\bm{\gamma}}_{1}(w)$ is that
$\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}0$ ( 3.2)
for any value of parameters satisfying underlying assumptions, where $\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}$ denotes convergence in probability. Note that (3.2) is equivalent to $\mathsf{E}[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|^{2}]\to 0$, because $0\leq\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|\leq\sqrt{2}$. Clearly, whether $\bm{\alpha}=\bm{0}_{q}$ or not is crucial for $\hat{\bm{\gamma}}_{1}(w)$ with $w\in[0,1)$, so, in some settings, consistency will be discussed under the assumption that $\bm{\alpha}\neq\bm{0}_{q}$.
Remark 6.
We use Lemma 3.1 to bound $\mathsf{E}[\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\|^{2}_{\rm op}]$ for the purpose of getting a specific bound that will be used to determine the value of $w$ below. Evaluating $\mathsf{E}[\|\bm{S}(w)-\mathsf{E}[\bm{S}(w)]\|_{\rm op}]$ directly may provide a better bound. When $w=1$ or when $\bm{\alpha}=\bm{0}_{q},w=1/2$, results such as Theorem 4 of Koltchinskii and Lounici (2017) or Theorem 4.7.1 of Vershynin (2018) can be used. For example, when $\bm{\alpha}=\bm{0}_{q}$, it holds that
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}^{(T)}-\bm{\gamma}_{1}\|^{2}\right]\leq\mathsf{E}\left[\sqrt{2}\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}^{(T)}-\bm{\gamma}_{1}\|\right]\leq\frac{C}{\lambda_{1}-\lambda_{2}}\max\left\{\sqrt{\frac{\lambda_{1}{\rm tr}(\bm{\Sigma})}{n-1}},\frac{{\rm tr}(\bm{\Sigma})}{n-1}\right\},$
where $C$ is a positive constant. To discuss consistency, this bound is better than the bound in Remark 5 especially when $\lambda_{1}/{\rm tr}(\bm{\Sigma})$ is small.
In practice, we should determine the value of the weight $w$ for the estimator $\hat{\bm{\gamma}}_{1}(w)$. Other than $w=0,0.5,1$, it may be reasonable to select the value of $w$ minimizing the upper bound, the right-hand side of (3.1), for the mean squared error. The value is provided in the following theorem.
Theorem 3.4.
The right-hand side of (3.1) attains the unique minimum at
$w^{*}=\frac{q\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}}{2q\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}+\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}\|\bm{X}\bm{\alpha}\|^{2}}.$ ( 3.3)
Proof.
For simplicity, we let
$a={\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2},\ \ \ b=\lambda_{1}+{\rm tr}(\bm{\Sigma}),\ \ \ c=\|\bm{X}\bm{\alpha}\|^{2},\ \ \ d=\lambda_{1}-\lambda_{2}.$
Then, (3.1) is rewritten as
$8\frac{\{q(1-2w)+(n-1)w^{2}\}a+2(1-w)^{2}bc}{\left[\{q+(n-1-2q)w\}d+(1-w)c\right]^{2}},$
and so the value of $w$ minimizing (3.1) is equivalent to that minimizing
$f(w)=\frac{\{q(1-2w)+(n-1)w^{2}\}a+2(1-w)^{2}bc}{2\left[\{q+(n-1-2q)w\}d+(1-w)c\right]^{2}}.$
Since $\{q+(n-1-2q)w\}d>0$ for $0\leq w\leq 1$ due to $q+(n-1-2q)w=q(1-w)+(n-1-q)w>0$ and $d=\lambda_{1}-\lambda_{2}>0$, the function $f(w)$ is continuous on $0\leq w\leq 1$. Differentiating $f(w)$ with respect to $w$, we have
$\frac{\mathrm{d}f(w)}{\mathrm{d}w}$ =$=\frac{\{-2q+2(n-1)w\}a-4(1-w)bc}{2\left[\{q+(n-1-2q)w\}d+(1-w)c\right]^{2}}-\frac{\left[\{q(1-2w)+(n-1)w^{2}\}a+2(1-w)^{2}bc\right]\left\{(n-1-2q)d-c\right\}}{\left[\{q+(n-1-2q)w\}d+(1-w)c\right]^{3}}$
=$=\frac{\left\{(2adq+ac+2bcd)w-(adq+2bcd)\right\}(n-1-q)}{\left[\{q+(n-1-2q)w\}d+(1-w)c\right]^{3}},$
from which we see that $f(w)$ is strictly decreasing on
$0\leq w<\frac{adq+2bcd}{2adq+2bcd+ac}$
and strictly increasing on
$\frac{adq+2bcd}{2adq+2bcd+ac}<w\leq 1.$
Hence, the value of $w$ minimizing the function $f(w)$ is given by
$w^{*}=\frac{adq+2bcd}{2adq+2bcd+ac}.$
Substituting $a,b,c,d$, we have (3.3). This completes the proof. ∎
A natural estimator for $w^{*}$ will be given in Section 6. We see from Theorem 3.4 that (i) if $\|\bm{X}\bm{\alpha}\|^{2}\approx 0$, then $w^{*}\approx 0.5$, (ii) if $\lambda_{1}-\lambda_{2}\approx 0$, then $w^{*}\approx 0$. Moreover, we have the bound $w^{*}<2/3$, because it follows from
$(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})<\lambda_{1}^{2}+\lambda_{1}{\rm tr}(\bm{\Sigma})<{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}$
that
$w^{*}$ =$=\frac{\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})q+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}}{2\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})q+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}+\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}\|\bm{X}\bm{\alpha}\|^{2}}$
<$<\frac{\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})q+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}}{2\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})q+3(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}}\leq\frac{2}{3}.$
4 Consistency under weak identifiability regimes
In this section, we discuss the case of $n\to\infty$ with fixed $p,q$. Assume that $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a positive-definite matrix as $n\to\infty$. Let $\bm{\alpha}$ be fixed. Then, $\|\bm{X}\bm{\alpha}\|^{2}=\Theta(n)$ holds for $\bm{\alpha}\neq\bm{0}_{q}$, and $\|\bm{X}\bm{\alpha}\|^{2}=0$ holds for $\bm{\alpha}=\bm{0}_{q}$, where $a_{n}=\Theta(n^{u})$ for a sequence $\{a_{n}\}_{n\in\mathbb{N}}$ means that lim $\liminf(a_{n}/n^{u})>0$ and lim $\limsup(a_{n}/n^{u})<\infty$ as $n\to\infty$. If $\bm{\Sigma}$ is fixed (with $\lambda_{1}-\lambda_{2}>0$), then it is obvious that $\hat{\bm{\gamma}}_{1}(w)$ is always a consistent estimator of $\bm{\gamma}_{1}$ up to sign for all $w\in[0,1]$, because the upper bound (3.1) for the mean squared error of $\hat{\bm{\gamma}}_{1}(w)$ converges to 00 as follows:
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|^{2}\right]$
$\leq\frac{8\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=\frac{8\{q(1-w)^{2}+(n-1-q)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q(1-w)+(n-1-q)w\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=\frac{\Theta(n)}{\left(\Theta(n)\right)^{2}}=\Theta\left(n^{-1}\right).$
Hence, in this section, we focus on the case of $\lambda_{1}-\lambda_{2}\to 0$ as $n\to\infty$. Such regimes are called weak identifiability regimes in Paindaveine et al. (2020) and Paindaveine and Verdebout (2023) in which some specific structures for $\bm{\Sigma}$ are assumed. These regimes correspond to the cases where the eigenvector of $\bm{\Sigma}$ corresponding to the largest eigenvalue is not identifiable in the limit. More specifically, we consider asymptotic regimes $n\to\infty$ under the following assumption.
Assumption 1.
As $n\to\infty$ with $q$, $p$, and $\bm{\alpha}$ held fixed, suppose the followings:
  • (i)
    $\bm{X}$ satisfies that $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a finite positive-definite matrix;
  • (ii)
    $\bm{\Sigma}$ satisfies that ${\rm tr}(\bm{\Sigma})=\Theta(1)$, ${\rm tr}(\bm{\Sigma}^{2})=\Theta(1)$, and $\lambda_{1}-\lambda_{2}=\Theta(n^{-\eta})$ for some $\eta>0$.
The consistency of $\hat{\bm{\gamma}}_{1}(w)$ up to sign is established in the following theorem.
Theorem 4.1.
Let Assumption 1 hold. Then, as $n\to\infty$ with $q$, $p$, and $\bm{\alpha}$ held fixed, sufficient conditions for (3.2) are given by
[…]
Proof.
First, we discuss the case of $\bm{\alpha}\neq\bm{0}_{q}$, that is, $\|\bm{X}\bm{\alpha}\|^{2}=\Theta(n)$. If $10<w<$ 1, then
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|^{2}\right]$
$\leq\frac{8\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=\frac{\Theta(n)+\Theta(n)}{\left(\Theta(n^{1-\eta})+\Theta(n)\right)^{2}}=\frac{\Theta(n)}{\left(\Theta(n)\right)^{2}}=\Theta\left(n^{-1}\right).$
If $w=0$, then
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(0)-\bm{\gamma}_{1}\|^{2}\right]$ $\leq\frac{8q\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left\{q(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right\}^{2}}$
$=\frac{\Theta(1)+\Theta(n)}{\left(\Theta(n^{-\eta})+\Theta(n)\right)^{2}}=\frac{\Theta(n)}{\left(\Theta(n)\right)^{2}}=\Theta\left(n^{-1}\right).$
If $w=1$, then
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(1)-\bm{\gamma}_{1}\|^{2}\right]$ $\leq\frac{8(n-1-q)\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}}{\left\{(n-1-q)(\lambda_{1}-\lambda_{2})\right\}^{2}}$
$=\frac{\Theta(n)}{\left(\Theta(n^{1-\eta})\right)^{2}}=\frac{\Theta(n)}{\Theta(n^{2-2\eta})}=\Theta\left(n^{2\eta-1}\right).$
Hence, when $\bm{\alpha}\neq\bm{0}_{q}$, (3.2) always holds for $10\leq w<$ 1 but (3.2) holds if $\eta<1/2$ for $w=1$.
Next, we discuss the case of $\bm{\alpha}=\bm{0}_{q}$. If $0<w\leq 1$, then
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|^{2}\right]$ $\leq\frac{8\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}}{\left[\{q+(n-1-2q)w\}(\lambda_{1}-\lambda_{2})\right]^{2}}$
$=\frac{\Theta(n)}{\left(\Theta(n^{1-\eta})\right)^{2}}=\frac{\Theta(n)}{\Theta(n^{2-2\eta})}=\Theta\left(n^{2\eta-1}\right).$
If $w=0$, then
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(0)-\bm{\gamma}_{1}\|^{2}\right]\leq\frac{8q\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}}{\left\{q(\lambda_{1}-\lambda_{2})\right\}^{2}}=\frac{\Theta(1)}{\left(\Theta(n^{-\eta})\right)^{2}}=\Theta\left(n^{2\eta}\right).$
Hence, when $\bm{\alpha}=\bm{0}_{q}$, (3.2) holds if $\eta<1/2$ for $0<w\leq 1$ but does not necessarily hold for $w=0$.
Summarizing the above results, we see that (3.2) with $10<w<$ 1 holds if at least one of $\bm{\alpha}\neq\bm{0}_{q}$ and $\eta<1/2$ is satisfied, (3.2) with $w=0$ holds if $\bm{\alpha}\neq\bm{0}_{q}$, and (3.2) with $w=1$ holds if $\eta<1/2$. ∎
Remark 7.
From Remarks 1 and 5, when $\bm{\alpha}=\bm{0}_{q}$, if $\lambda_{1}-\lambda_{2}\to 0$ and $n^{1/2}(\lambda_{1}-\lambda_{2})\to\infty$, then the leading principal eigenvector $\hat{\bm{\gamma}}_{1}^{(T)}$ is consistent. It is consistent with a result in Paindaveine et al. (2020, Lemma 2.3). Also from Paindaveine et al. (2020, Lemma 2.3), if $\limsup_{n\to\infty}n^{1/2}(\lambda_{1}-\lambda_{2})<\infty$ then $\hat{\bm{\gamma}}_{1}^{(T)}$ is no more consistent, because their setting is a special case of ours.
Next, the consistency of $\hat{\bm{\gamma}}_{1}(w^{*})$ up to sign is established as follows.
Theorem 4.2.
Let Assumption 1 hold. Then, as $n\to\infty$ with $q$, $p$, and $\bm{\alpha}$ held fixed, a sufficient condition for
$\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w^{*})-\bm{\gamma}_{1}\|\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}0$ ( 4.1)
is given by
[…]
Proof.
We notice that $w^{*}=1/2$ under $\bm{\alpha}=\bm{0}_{q}$ and $w^{*}=\Theta(n^{-\eta})$ under $\bm{\alpha}\neq\bm{0}_{q}$. For $\bm{\alpha}=\bm{0}_{q}$, since $w^{*}=1/2$, we see from Theorem 4.1 that $\eta<1/2$ is a sufficient condition for (4.1). For $\bm{\alpha}\neq\bm{0}_{q}$,
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w^{*})-\bm{\gamma}_{1}\|^{2}\right]$
$\leq\frac{8\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q(1-w^{*})+(n-1-q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=\frac{\Theta(1)+\Theta(n^{1-2\eta})+\Theta(n)}{\left(\Theta(n^{-\eta})+\Theta(n^{1-2\eta})+\Theta(n)\right)^{2}}=\frac{\Theta(n)}{\left(\Theta(n)\right)^{2}}=\Theta\left(n^{-1}\right),$
which ensures that (4.1) holds. ∎
5 Consistency under large $p$, large $n$ regimes
In this section, we consider “large $p$, large $n$” regimes under which $n$ and $p$ tend to positive infinity simultaneously. These regimes are frequently discussed in the literature. We suppose the following assumption.
Assumption 2.
As $n$ and $p$ tend to $\infty$ satisfying that ${n}/{p^{\delta}}$ converges to a finite positive value for some $\delta\in(0,\infty)$ with $q$ held fixed, suppose the follwings:
  • (i)
    $\bm{X}$ and $\bm{\alpha}$ satisfy that ${\|\bm{X}\bm{\alpha}\|^{2}}=\Theta(p^{\delta})$;
  • (ii)
    $\bm{\Sigma}$ satisfies that ${{\rm tr}(\bm{\Sigma})}=\Theta(p)$ and $(\lambda_{1}-\lambda_{2})=\Theta(p^{\beta})$ for some $\beta\in(-\infty,1]$.
Under Assumption 2, $\beta>0$ implies that the largest eigenvalue is (asymptotically) spiked, while non-spiked eigenvalues imply $\beta\leq 0$. The cases of $\beta<0$ correspond to weak identifiability regimes. Moreover, $\delta<1$ implies $p/n\to\infty$.
Remark 8.
Spiked eigenvalue assumptions on covariance matrices are typically assumed in high-dimensional multivariate analysis; see Shen et al. (2016), Wang and Fan (2017), Yata and Aoshima (2013), and so on.
The following theorem provides sufficient conditions for the consistency up to sign of $\hat{\bm{\gamma}}_{1}(w)$. Recall that the meaning of the consistency is explained in (3.2) that works even when the dimension $p$ of the parameter $\bm{\gamma}_{1}$ diverges.
Theorem 5.1.
Let Assumption 2 hold. Then, as $p$ and $n$ tend to $\infty$ satisfying that ${n}/{p^{\delta}}$ converges to a finite positive value with $q$ held fixed, sufficient conditions for (3.2) are given by
$\begin{array}[]{ll}\delta>1&\mbox{for $w=0$},\\ \delta>\min\{2-2\beta,2\}&\mbox{for $0<w<1$},\\ \delta>2-2\beta&\mbox{for $w=1$}.\end{array}$
Proof.
For $10<w<$ 1, it holds that
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(w)-\bm{\gamma}_{1}\|^{2}\right]$
$\leq\frac{8\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=\frac{\Theta(p^{2+\delta})+\Theta(p^{1+\delta})}{\left(\Theta(p^{\beta+\delta})+\Theta(p^{\delta})\right)^{2}}=\frac{\Theta(p^{2+\delta})}{\Theta(p^{\max\{2\beta+2\delta,2\delta\}})}=\Theta\left(p^{\min\{2-2\beta-\delta,2-\delta\}}\right),$
which means that if $\delta>\min\{2-2\beta,2\}$, then the mean squared error of $\hat{\bm{\gamma}}_{1}(w)$ converges to 00, and thus (3.2) holds.
For $w=1$,
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(1)-\bm{\gamma}_{1}\|^{2}\right]\leq\frac{8\{n-1-q\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}}{(n-1-q)^{2}(\lambda_{1}-\lambda_{2})^{2}}=\frac{\Theta(p^{2+\delta})}{\Theta(p^{2\beta+2\delta})}=\Theta\left(p^{2-2\beta-\delta}\right),$
from which we see that if $\delta>2-2\beta$, then (3.2) holds.
For $w=0$,
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(0)-\bm{\gamma}_{1}\|^{2}\right]$
$\leq\frac{8q\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left\{q(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right\}^{2}}$
$=\frac{\Theta(p^{2})+\Theta(p^{1+\delta})}{\left(\Theta(p^{\beta})+\Theta(p^{\delta})\right)^{2}}=\frac{\Theta(p^{\max\{2,1+\delta\}})}{\Theta(p^{\max\{2\beta,2\delta\}})}=\left\{\begin{array}[]{ll}\Theta\left(p^{1-\delta}\right)&(\delta\geq 1),\\ \Theta\left(p^{\min\{2-2\beta,2-2\delta\}}\right)&(\delta<1).\end{array}\right.$
Since $\min\{2-2\beta,2-2\delta\}\geq 0$ holds for $\delta<1$, we see that (3.2) holds for $\delta>1$.
This completes the proof. ∎
Theorem 5.1 implies that sufficient conditions for the consistency of $\hat{\bm{\gamma}}_{1}(w)$ with $10<w<$ 1 is (i) weaker than $\hat{\bm{\gamma}}_{1}(1)$ for $\beta<0$ (weak identifiability regimes) and the same as $\hat{\bm{\gamma}}_{1}(1)$ otherwise, (ii) weaker than $\hat{\bm{\gamma}}_{1}(0)$ for $\beta>0.5$ (strongly spiked eigenvalue regimes) but stronger than $\hat{\bm{\gamma}}_{1}(0)$ for $\beta<0.5$ (weakly spiked eigenvalue or weak identifiability regimes). We see that if $\beta>0.5$ (strongly spiked eigenvalue regimes), then the consistency of $\hat{\bm{\gamma}}_{1}(w)$ with $10<w<$ 1 remains valid even for ${p}/{n}\to\infty$ as long as ${p^{2-2\beta}}/{n}\to 0$ holds. We also see that, even for $\beta<0$ (weak identifiability regimes), we have the consistency of $\hat{\bm{\gamma}}_{1}(w)$ with $10<w<$ 1 if ${p^{2}}/{n}\to 0$.
Next theorem provides a sufficient condition for the consistency up to sign of $\hat{\bm{\gamma}}_{1}(w^{*})$.
Theorem 5.2.
Let Assumption 2 hold. Then, as $p$ and $n$ tend to $\infty$ satisfying that ${n}/{p^{\delta}}$ converges to a finite positive value with $q$ held fixed, a sufficient condition for (4.1) is given by
$\delta>\min\{2-2\beta,1\}.$
Proof.
Since
$w^{*}$ =$=\frac{\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})q+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}}{2\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})q+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}+\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}\|\bm{X}\bm{\alpha}\|^{2}}$
$=\frac{\Theta(p^{2+\beta})+\Theta_{p}(p^{1+\beta+\delta})}{\Theta_{p}(p^{2+\beta})+\Theta_{p}(p^{1+\beta+\delta})+\Theta_{p}(p^{2+\delta})}=\frac{\Theta(p^{\max\{2+\beta,1+\beta+\delta\}})}{\Theta_{p}(p^{\max\{2+\beta,2+\delta\}})},$
we notice that, for $\delta\leq 1$, $w^{*}=\Theta(1)$ under $\beta\geq\delta$ and $w^{*}=\Theta(p^{\beta-\delta})$ under $\beta<\delta$. For $\delta>1$, $w^{*}=\Theta(p^{\beta-1})$. From $30\leq w^{*}<2/$ 3, it follows that $1-w^{*}=\Theta(1)$. We also note that
$\frac{8\{q(1-2w^{*})+(n-1)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=\frac{8\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q(1-w^{*})+(n-1-q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}.$
First, we discuss the case of $\delta>1$. It holds that
$\frac{8\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q(1-w^{*})+(n-1-q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=\frac{\Theta(p^{\max\{0,\delta+2\beta-2\}})\Theta(p^{2})+\Theta(p^{\delta+1})}{\left(\Theta(p^{\max\{0,\delta+\beta-1\}})\Theta(p^{\beta})+\Theta(p^{\delta})\right)^{2}}=\frac{\Theta(p^{\max\{2,\delta+1,\delta+2\beta\}})}{\left(\Theta(p^{\max\{\delta,\beta,\delta+2\beta-1\}})\right)^{2}}$
$=\frac{\Theta(p^{\max\{\delta+1,\delta+2\beta\}})}{\left(\Theta(p^{\max\{\delta,\delta+2\beta-1\}})\right)^{2}}.$
If $\beta>1/2$, then
$\frac{\Theta(p^{\max\{\delta+1,\delta+2\beta\}})}{\left(\Theta(p^{\max\{\delta,\delta+2\beta-1\}})\right)^{2}}=\frac{\Theta(p^{\delta+2\beta})}{\left(\Theta(p^{\delta+2\beta-1})\right)^{2}}=\frac{\Theta(p^{\delta+2\beta})}{\Theta(p^{2\delta+4\beta-2})}=\Theta(p^{-\delta-2\beta+2}),$
which implies (4.1) because $\delta>1$ and $\beta>1/2$ ensure $-\delta-2\beta+2<0$. If $\beta\leq 1/2$, then
$\frac{\Theta(p^{\max\{\delta+1,\delta+2\beta\}})}{\left(\Theta(p^{\max\{\delta,\delta+2\beta-1\}})\right)^{2}}=\frac{\Theta(p^{\delta+1})}{\left(\Theta(p^{\delta})\right)^{2}}=\frac{\Theta(p^{\delta+1})}{\Theta(p^{2\delta})}=\Theta(p^{-\delta+1}),$
which also implies (4.1).
Next, we deal with the case of $\delta\leq 1$ and $\beta<\delta$. It holds that
$\frac{8\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q(1-w^{*})+(n-1-q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=\frac{\Theta(p^{\max\{0,-\delta+2\beta\}})\Theta(p^{2})+\Theta(p^{\delta+1})}{\left(\Theta(p^{\max\{0,\beta\}})\Theta(p^{\beta})+\Theta(p^{\delta})\right)^{2}}=\frac{\Theta(p^{\max\{2,-\delta+2\beta+2\}})+\Theta(p^{\delta+1})}{\left(\Theta(p^{\max\{\beta,2\beta\}})+\Theta(p^{\delta})\right)^{2}}$
$=\frac{\Theta(p^{\max\{2,-\delta+2\beta+2\}})}{\left(\Theta(p^{\max\{\delta,2\beta\}})\right)^{2}}.$
If $\delta>2\beta$, then
$\frac{\Theta(p^{\max\{2,-\delta+2\beta+2\}})}{\left(\Theta(p^{\max\{\delta,2\beta\}})\right)^{2}}=\frac{\Theta(p^{2})}{\left(\Theta(p^{\delta})\right)^{2}}=\frac{\Theta(p^{2})}{\Theta(p^{2\delta})}=\Theta(p^{-2\delta+2}),$
from which we see that (4.1) does not necessarily hold in this case. On the other hand, if $\beta<\delta<2\beta$, then
$\frac{\Theta(p^{\max\{2,-\delta+2\beta+2\}})}{\left(\Theta(p^{\max\{\delta,2\beta\}})\right)^{2}}=\frac{\Theta(p^{-\delta+2\beta+2})}{\left(\Theta(p^{2\beta})\right)^{2}}=\frac{\Theta(p^{-\delta+2\beta+2})}{\Theta(p^{4\beta})}=\Theta(p^{-\delta-2\beta+2}),$
which indicates that (4.1) holds if $\delta>2-2\beta$.
Finally, we discuss the case of $\delta\leq 1$ and $\beta\geq\delta$. It holds that
$\frac{8\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left(\{q(1-w^{*})+(n-1-q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right)^{2}}$
$=\frac{\Theta(p^{\delta})\Theta(p^{2})+\Theta(p^{\delta+1})}{\left(\Theta(p^{\delta})\Theta(p^{\beta})+\Theta(p^{\delta})\right)^{2}}=\frac{\Theta(p^{\delta+2})}{\left(\Theta(p^{\delta+\beta})\right)^{2}}=\frac{\Theta(p^{\delta+2})}{\Theta(p^{2\delta+2\beta})}=\Theta(p^{-\delta-2\beta+2}),$
from which we see that (4.1) holds if $\delta>2-2\beta$.
From what has already been proved, noting that $2\beta>2-\delta$ is not satisfied under $\delta\leq 1$ and $\delta>2\beta$ ($2\beta<\delta\leq 2-\delta$), we conclude that (4.1) always holds under $\delta>1$ and that (4.1) holds for $\delta>2-2\beta$ under $\delta\leq 1$. ∎
We see that, as for sufficient conditions for the consistency, $\hat{\bm{\gamma}}_{1}(w^{*})$ has the good points both of $\hat{\bm{\gamma}}_{1}(w)$ with $10<w<$ 1 and $\hat{\bm{\gamma}}_{1}(0)$. This may be because the inequality $\|\cdot\|_{\mathrm{op}}\leq\|\cdot\|_{\mathrm{F}}$ is used. We may relax sufficient conditions in Theorem 5.1 by avoiding this inequality.
6 Feasible $\hat{\bm{\gamma}}_{1}(w^{*})$ estimator
Since $w^{*}$ includes unknown parameters, in order to use $\hat{\bm{\gamma}}_{1}(w^{*})$, $w^{*}$ needs to be estimated. In this section, $n>2+q$ is additionally assumed. Then, a natural estimator of $w^{*}$ is given by
$\hat{w}^{*}=\frac{q\{\widehat{{\rm tr}(\bm{\Sigma}^{2})}+({\rm tr}(\hat{\bm{\Sigma}}))^{2}\}(\hat{\lambda}_{1}-\hat{\lambda}_{2})+2(\hat{\lambda}_{1}+{\rm tr}(\hat{\bm{\Sigma}}))(\hat{\lambda}_{1}-\hat{\lambda}_{2})\widehat{\|\bm{X}\bm{\alpha}\|^{2}}}{2q\{\widehat{{\rm tr}(\bm{\Sigma}^{2})}+({\rm tr}(\hat{\bm{\Sigma}}))^{2}\}(\hat{\lambda}_{1}-\hat{\lambda}_{2})+2(\hat{\lambda}_{1}+{\rm tr}(\hat{\bm{\Sigma}}))(\hat{\lambda}_{1}-\hat{\lambda}_{2})\widehat{\|\bm{X}\bm{\alpha}\|^{2}}+\{\widehat{{\rm tr}(\bm{\Sigma}^{2})}+({\rm tr}(\hat{\bm{\Sigma}}))^{2}\}\widehat{\|\bm{X}\bm{\alpha}\|^{2}}},$
where
$\hat{\bm{\Sigma}}=\frac{\bm{S}_{E}}{n-1-q},$
$\widehat{{\rm tr}(\bm{\Sigma}^{2})}=\frac{1}{(n+1-q)(n-2-q)}\left\{{\rm tr}(\bm{S}_{E}^{2})-\frac{\left({\rm tr}(\bm{S}_{E})\right)^{2}}{n-1-q}\right\},$
$\widehat{\|\bm{X}\bm{\alpha}\|^{2}}={\rm tr}(\bm{S}_{R})-\frac{q}{n-1-q}{\rm tr}(\bm{S}_{E}),$
and $\hat{\lambda}_{1}$ and $\hat{\lambda}_{2}$ are the largest and second largest eigenvalues of $\hat{\bm{\Sigma}}$, respectively. It holds that $\hat{w}^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}w^{*}$ under some asymptotic regimes such as a traditional large sample regime where $n\to\infty$ with the assumption that $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a positive-definite matrix while other parameters are fixed. Neither the consistency of $\hat{\bm{\gamma}}_{1}(w^{*})$ nor $\hat{w}^{*}$ necessarily means that a feasible estimator $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is consistent up to sign, that is to say,
$\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(\hat{w}^{*})-\bm{\gamma}_{1}\|\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}0.$ ( 6.1)
Hence, let us discuss asymptotic regimes under which $\hat{w}^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}w^{*}$ implies (6.1).
Theorem 6.1.
Under asymptotic regimes satisfying
$\frac{\{q(1-2w^{*})+(n-1)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}\to 0$ ( 6.2)
and
$\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)({\rm tr}(\bm{\Sigma}))^{2}-(n-2q)\lambda_{1}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}=O(1),$ ( 6.3)
if $\hat{w}^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}w^{*}$, then (6.1) holds.
Proof.
Since $\bm{\gamma}_{1}$ is the first principal eigenvector of $\mathsf{E}[\bm{S}(w^{*})]$ and since
$\|\bm{S}(\hat{w}^{*})-\mathsf{E}[\bm{S}(w)]\|_{\rm op}$ $\leq\|\bm{S}(\hat{w}^{*})-\bm{S}(w^{*})\|_{\rm op}+\|\bm{S}(w^{*})-\mathsf{E}[\bm{S}(w^{*})]\|_{\rm op}$
$=|\hat{w}^{*}-w^{*}|\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm op}+\|\bm{S}(w^{*})-\mathsf{E}[\bm{S}(w^{*})]\|_{\rm op},$
we have, by using the Davis–Kahan theorem,
$\sin\angle(\bm{\gamma}_{1},\hat{\bm{\gamma}}_{1}(\hat{w}^{*}))\leq\frac{2\|\bm{S}(\hat{w}^{*})-\mathsf{E}[\bm{S}(w^{*})]\|_{{\rm op}}}{\lambda_{1}(w^{*})-\lambda_{2}(w^{*})}\leq 2|\hat{w}^{*}-w^{*}|\frac{\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm op}}{\lambda_{1}(w^{*})-\lambda_{2}(w^{*})}+2\frac{\|\bm{S}(w^{*})-\mathsf{E}[\bm{S}(w^{*})]\|_{\rm op}}{\lambda_{1}(w^{*})-\lambda_{2}(w^{*})}.$
From (6.2), the second term on the right-hand side converges to 0 in probability. Thus, it suffices to show that
$\frac{\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm op}]}{\lambda_{1}(w^{*})-\lambda_{2}(w^{*})}=O(1)$
because $\hat{w}^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}w^{*}$ is assumed. Since the Jensen inequality provides $\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm op}]\leq\sqrt{\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm op}^{2}]}$, we evaluate $\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|^{2}_{\rm op}]$. It follows from
$\|\bm{S}_{R}-\bm{S}_{E}\|_{\mathrm{op}}^{2}$ $\leq\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm F}^{2}={\rm tr}\Bigl{(}(\bm{S}_{R}-\bm{S}_{E})(\bm{S}_{R}-\bm{S}_{E})^{\top}\Bigr{)}$
$={\rm tr}\Bigl{(}(\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}])^{2}\Bigr{)}+{\rm tr}\Bigl{(}(\bm{S}_{E}-\mathsf{E}[\bm{S}_{E}])^{2}\Bigr{)}+{\rm tr}\Bigl{(}(\mathsf{E}[\bm{S}_{R}-\bm{S}_{E}])^{2}\Bigr{)}$
$\quad-2{\rm tr}\Bigl{(}(\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}])(\bm{S}_{E}-\mathsf{E}[\bm{S}_{E}])\Bigr{)}-2{\rm tr}\Bigl{(}(\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}])(E[\bm{S}_{R}-\bm{S}_{E}])\Bigr{)}$
$\quad-2{\rm tr}\Bigl{(}(\bm{S}_{E}-\mathsf{E}[\bm{S}_{E}])(E[\bm{S}_{R}-\bm{S}_{E}])\Bigr{)},$
that
$\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|_{\mathrm{op}}^{2}]\leq\mathsf{E}\left[{\rm tr}\left((\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}])^{2}\right)\right]+\mathsf{E}\left[{\rm tr}\left((\bm{S}_{E}-\mathsf{E}[\bm{S}_{E}])^{2}\right)\right]+{\rm tr}\left((\mathsf{E}[\bm{S}_{R}-\bm{S}_{E}])^{2}\right).$
Here, recall that
$\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{R}-E[\bm{S}_{R}]\right)^{2}\right)\right]=q{\rm tr}(\bm{\Sigma}^{2})+q\left({\rm tr}(\bm{\Sigma})\right)^{2}+2\left(\lambda_{1}+{\rm tr}(\bm{\Sigma})\right)\|\bm{X}\bm{\alpha}\|^{2},$
$\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{E}-E[\bm{S}_{E}]\right)^{2}\right)\right]=(n-1-q){\rm tr}(\bm{\Sigma}^{2})+(n-1-q)\left({\rm tr}(\bm{\Sigma})\right)^{2}.$
Moreover, it holds that
${\rm tr}\left((\mathsf{E}[\bm{S}_{R}-\bm{S}_{E}])^{2}\right)$ $={\rm tr}\left(\left\{-(n-1-2q)\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}+\|\bm{X}\bm{\alpha}\|^{2}\bm{\gamma}_{1}\bm{\gamma}_{1}^{\prime}\right\}^{2}\right)$
$=\left(n-1-2q\right)^{2}{\rm tr}(\bm{\Sigma}^{2})+\|\bm{X}\bm{\alpha}\|^{4}-2(n-1-2q)\lambda_{1}\|\bm{X}\bm{\alpha}\|^{2}.$
Hence, we have
$\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm op}]$
$\leq\sqrt{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)({\rm tr}(\bm{\Sigma}))^{2}+\|\bm{X}\bm{\alpha}\|^{4}+2\{{\rm tr}(\bm{\Sigma})-(n-2-2q)\lambda_{1}\}\|\bm{X}\bm{\alpha}\|^{2}}.$
Noting that $w^{*}<2/3$, $\lambda_{1}(w^{*})-\lambda_{2}(w^{*})=\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}$, and
$\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)\{{\rm tr}(\bm{\Sigma})\}^{2}+\|\bm{X}\bm{\alpha}\|^{4}+\{{\rm tr}(\bm{\Sigma})-(n-2-2q)\lambda_{1}\}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
=$=\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)\{{\rm tr}(\bm{\Sigma})\}^{2}+\{{\rm tr}(\bm{\Sigma})-(n-2-2q)\lambda_{1}\}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}+O(1)$
=$=\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)\{{\rm tr}(\bm{\Sigma})\}^{2}-(n-2q)\lambda_{1}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}+o(1)+O(1),$
we have the conclusion. ∎
Remark 9.
The estimator $\hat{w}^{*}$ in Theorem 6.1 can be replaced with arbitrary consistent estimator of $w^{*}$.
Remark 10.
The condition (6.3) can be improved by determining the rate of convergence of $\hat{w}^{*}$: under asymptotic regimes satisfying (6.2) and
$\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)({\rm tr}(\bm{\Sigma}))^{2}+\{{\rm tr}(\bm{\Sigma})-(n-2q)\lambda_{1}\}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}=o(r_{n}^{2})$
for some non-decreasing sequence $\{r_{n}\}_{n\in\mathbb{N}}$ of positive values $0<r_{1}\leq r_{2}\leq\cdots$ satisfying that $r_{n}(\hat{w}^{*}-w^{*})$ is uniformly tight, the same conclusion follows.
Under the traditional large sample regime, (6.2) and (6.3) are satisfied, so we have the following corollary.
Corollary 6.2.
Suppose that, as $n\to\infty$, $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a positive-definite matrix with other parameters held fixed. Then, as $n\to\infty$, (6.1) holds.
The performance of $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$, compared with $\hat{\bm{\gamma}}_{1}(w^{*})$, is examined through numerical simulations in Section 7.
7 Numerical simulations
This section provides the results of some numerical simulations for comparing performances of the estimators $\hat{\bm{\gamma}}_{1}^{(T)}=\hat{\bm{\gamma}}_{1}(0.5)$, $\hat{\bm{\gamma}}_{1}^{(E)}=\hat{\bm{\gamma}}_{1}(1)$, $\hat{\bm{\gamma}}_{1}^{(R)}=\hat{\bm{\gamma}}_{1}(0)$, $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$, and $\hat{\bm{\gamma}}_{1}(w^{*})$. For comparison, $\hat{\bm{\gamma}}_{1}(w)$ with $w=0.1,0.2,0.3,0.4,0.6$ are also included. The estimators $\hat{\bm{\gamma}}_{1}(w)$ with $w=0.7,0.8,0.9$ are not included since $w^{*}<2/3$ is shown in Section 3.
First, we consider a traditional large sample regime and a weak identifiability regime, as follows. The number of response variables $p$ is 10, the number of explanatory variables $q$ is 5, and the sample size $n$ is $20,50,100,200,500$. The explanatory variables matrix $\bm{X}$ ($n\times q$ matrix) is a random sample of size $n$ drawn from $\mathcal{N}_{q}(\bm{0}_{q},\bm{I}_{q})$ with centralization ($\bm{X}^{\prime}\bm{1}_{n}=\bm{0}_{q}$). Let $\bm{\alpha}=\bm{1}_{q}$ (the $q\times 1$ vector whose all elements are $1$). Let $\bm{\Sigma}=\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}$, where we set the diagonal matrix of eigenvalues $\bm{\Lambda}={\rm diag}(\lambda_{1},\lambda_{2},\dots,\lambda_{p})$ by $\lambda_{1}=2$ (traditional large sample regime) or $\lambda_{1}=1+1/n$ (weak identifiability regime) and by $\lambda_{i}=1,\ i=2,3,\dots,p$ and we also set the orthogonal matrix of eigenvectors $\bm{\Gamma}=(\bm{\gamma}_{1},\bm{\gamma}_{2},\dots,\bm{\gamma}_{p})$ by using normalized eigenvectors of a sample covariance matrix of the random sample of size $2p$ drawn from $\mathcal{N}_{p}(\bm{0}_{p},\bm{I}_{p})$. We note that the setting of $\bm{\Gamma}$ does not affect the results of simulations. Table 1-(a) (traditional large sample regime) and (b) (weak identifiability regime) show the results of mean squared errors of the estimators based on 1000 numerical simulations.
We see from Table 1-(a) that $\hat{\bm{\gamma}}_{1}(0.3)$ is the best for all cases and that $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ and $\hat{\bm{\gamma}}_{1}(w^{*})$ are better than both $\hat{\bm{\gamma}}_{1}^{(R)}$ and $\hat{\bm{\gamma}}_{1}^{(T)}$. These results suggest that the optimal value of $w$, which is unknown in practice, is in $(0,1/2)$ and motivate to use the proposed estimator $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$. As previously anticipated, $\hat{\bm{\gamma}}_{1}^{(E)}$ is not better than other estimators.
We see from Table 1-(b) that $\hat{\bm{\gamma}}_{1}^{(R)}$ and $\hat{\bm{\gamma}}_{1}(w^{*})$ are the best for almost all cases. As an exception, for $n=20$, $\hat{\bm{\gamma}}_{1}(0.1)$ is the best, and $\hat{\bm{\gamma}}_{1}(w^{*})$ is slightly better than $\hat{\bm{\gamma}}_{1}^{(R)}$. As for $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$, it is not good as $\hat{\bm{\gamma}}_{1}(w^{*})$ for small $n$, but their difference goes to 00 as $n$ increases. In contrast, $\hat{\bm{\gamma}}_{1}^{(E)}$ is quite bad, which may be due to the fact from Theorem 4.1 that $\hat{\bm{\gamma}}_{1}^{(E)}$ is not necessarily consistent because $\eta=1$ does not satisfy $\eta<1/2$. The inconsistency of $\hat{\bm{\gamma}}_{1}^{(E)}$ may affect the performance of $\hat{\bm{\gamma}}_{1}^{(T)}$, which is worse than $\hat{\bm{\gamma}}_{1}^{(R)}$, $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$, and $\hat{\bm{\gamma}}_{1}(w^{*})$.
[…]
We see from Table 1-(c) that $\hat{\bm{\gamma}}_{1}^{(T)}$ is the best for all cases. Although $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ and $\hat{\bm{\gamma}}_{1}(w^{*})$ are not the best, their performances become closer to $\hat{\bm{\gamma}}_{1}^{(T)}$ as $p$ and $n$ increase. As for $\hat{\bm{\gamma}}_{1}^{(E)}$, its performance is bad for small $p$ and rapidly closer to $\hat{\bm{\gamma}}_{1}^{(T)}$ as $p$ and $n$ increase, which may be due to the fact from Theorem 5.1 that $\hat{\bm{\gamma}}_{1}^{(E)}$ is consistent because $\delta=0.8$ and $\beta=0.8$ satisfy $\delta>2-2\beta$. In contrast, $\hat{\bm{\gamma}}_{1}^{(R)}$ is bad especially for larger $p$, which may be due to the fact from Theorem 5 that $\hat{\bm{\gamma}}_{1}^{(R)}$ is not necessarily consistent because $\delta=0.8$ does not satisfy $\delta>1$.
In our settings, $\hat{\bm{\gamma}}_{1}^{(T)}$ and $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ are good in all cases examined. In particular, on the one hand $\hat{\bm{\gamma}}_{1}^{(T)}$ is better in typical settings; on the other hand $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is better in settings where $\lambda_{1}-\lambda_{2}$ is small.
Table 1: Numerical results of the mean squared errors of $\hat{\bm{\gamma}}_{1}^{(T)}$, $\hat{\bm{\gamma}}_{1}^{(E)}$, $\hat{\bm{\gamma}}_{1}^{(R)}$, $\hat{\bm{\gamma}}_{1}(w)$ with $w=0.1,0.2,0.3,0.4,0.6$, $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ (with the average of the values of $\hat{w}^{*}$), and $\hat{\bm{\gamma}}_{1}(w^{*})$ (with the average of $w^{*}$) based on 1000 numerical simulations. The best result up to significant digits in each setting is in bold face.
(a) Results in the case under a traditional large sample regime.
$n$ $\hat{\bm{\gamma}}_{1}^{(T)}$ $\hat{\bm{\gamma}}_{1}^{(E)}$ $\hat{\bm{\gamma}}_{1}^{(R)}$ $\hat{\bm{\gamma}}_{1}(0.1)$ $\hat{\bm{\gamma}}_{1}(0.2)$ $\hat{\bm{\gamma}}_{1}(0.3)$ $\hat{\bm{\gamma}}_{1}(0.4)$ $\hat{\bm{\gamma}}_{1}(0.6)$ $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ $\hat{w}^{*}$ $\hat{\bm{\gamma}}_{1}(w^{*})$ $w^{*}$ 20 0.10517 0.90481 0.10723 0.10404 0.10113 0.09915 0.09949 0.12335 0.10101 (0.19359) 0.10109 (0.19237) 50 0.03734 0.46512 0.03887 0.03740 0.03613 0.03531 0.03540 0.04309 0.03684 (0.12519) 0.03642 (0.17417) 100 0.01831 0.22117 0.01873 0.01804 0.01745 0.01711 0.01724 0.02127 0.01782 (0.11918) 0.01762 (0.16823) 200 0.00862 0.09505 0.00886 0.00852 0.00823 0.00806 0.00812 0.01000 0.00843 (0.12590) 0.00832 (0.16535) 500 0.00349 0.03515 0.00363 0.00349 0.00337 0.00330 0.00331 0.00400 0.00344 (0.13702) 0.00341 (0.16363)
(b) Results in the case under a weak identifiability regime $\lambda_{1}-\lambda_{2}=1/n$.
$n$ $\hat{\bm{\gamma}}_{1}^{(T)}$ $\hat{\bm{\gamma}}_{1}^{(E)}$ $\hat{\bm{\gamma}}_{1}^{(R)}$ $\hat{\bm{\gamma}}_{1}(0.1)$ $\hat{\bm{\gamma}}_{1}(0.2)$ $\hat{\bm{\gamma}}_{1}(0.3)$ $\hat{\bm{\gamma}}_{1}(0.4)$ $\hat{\bm{\gamma}}_{1}(0.6)$ $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ $\hat{w}^{*}$ $\hat{\bm{\gamma}}_{1}(w^{*})$ $w^{*}$ 20 0.12627 1.44891 0.10464 0.10456 0.10508 0.10699 0.11307 0.16196 0.10721 (0.17238) 0.10461 (0.01271) 50 0.04431 1.47012 0.03730 0.03733 0.03762 0.03841 0.04022 0.05414 0.03733 (0.07300) 0.03730 (0.00440) 100 0.02267 1.46049 0.01857 0.01865 0.01888 0.01939 0.02044 0.02773 0.01860 (0.04313) 0.01857 (0.00210) 200 0.01078 1.47838 0.00905 0.00906 0.00914 0.00934 0.00979 0.01306 0.00905 (0.02688) 0.00905 (0.00102) 500 0.00442 1.48953 0.00365 0.00366 0.00371 0.00380 0.00400 0.00536 0.00365 (0.01564) 0.00365 (0.00040)
(c) Results in the case under a large $p$, large $n$ regime $n=\lfloor p^{0.8}\rfloor$, $\lambda_{1}=p^{0.8}$, $\lambda_{2}=p^{0.4}$.
$p$ $\hat{\bm{\gamma}}_{1}^{(T)}$ $\hat{\bm{\gamma}}_{1}^{(E)}$ $\hat{\bm{\gamma}}_{1}^{(R)}$ $\hat{\bm{\gamma}}_{1}(0.1)$ $\hat{\bm{\gamma}}_{1}(0.2)$ $\hat{\bm{\gamma}}_{1}(0.3)$ $\hat{\bm{\gamma}}_{1}(0.4)$ $\hat{\bm{\gamma}}_{1}(0.6)$ $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ $\hat{w}^{*}$ $\hat{\bm{\gamma}}_{1}(w^{*})$ $w^{*}$ 20 0.21682 0.63574 0.32229 0.29480 0.26881 0.24459 0.22383 0.23027 0.22046 (0.43745) 0.21935 (0.43548) 50 0.10300 0.16489 0.29513 0.21403 0.15606 0.12276 0.10712 0.10690 0.10730 (0.43157) 0.10457 (0.43582) 100 0.06570 0.08581 0.30129 0.16516 0.10299 0.07761 0.06789 0.06735 0.06790 (0.43595) 0.06662 (0.43235) 200 0.04191 0.04842 0.32524 0.11736 0.06423 0.04807 0.04296 0.04251 0.04286 (0.44012) 0.04241 (0.42782) 500 0.02412 0.02587 0.35335 0.06643 0.03371 0.02641 0.02447 0.02430 0.02451 (0.43357) 0.02433 (0.42058)
Acknowledgments
This study was supported in part by Japan Society for the Promotion of Science KAKENHI Grant Numbers 20K11713 and 21K13836.